For nearly 20 years, Custom Coatings has provided painting and galvanizing services for manufacturers in its region. Manufacturers of various metal products have relied on the quality and quick turnaround time provided by Custom Coatings and its 20 skilled employees. During the last year, as a result of a sharp upturn in the economy, the company’s sales have increased by 30% relative to the previous year. The company has not been able to increase its capacity fast enough, so Custom Coatings has had to turn work away because it cannot keep up with customer requests.

Top management is considering the purchase of a sophisticated robotic painting booth. The booth would represent a considerable move in the direction of automation versus manual labour. If Custom Coatings purchases the booth, it would most likely lay off 15 of its skilled painters. To analyze the decision, the company compiled production information from the most recent year and then prepared a parallel compilation assuming that the company would purchase the new equipment and lay off the workers. The data are shown at right. As you can see, the company projects that during the last year it would have been far more profitable if it had used the automated approach.
Current Approach Automated Approach
Sales $2,000,000 $2,000,000
Variable costs 1,200,000 400,000
Contribution margin 800,000 1,600,000
Fixed costs 200,000 600,000
Operating income $600,000 $1,000,000
Calculate the contribution margin ratio under each approach.
Current Approach Automated Approach
Contribution margin ratio

%

%

LINK TO TEXT

LINK TO TEXT

LINK TO TEXT

LINK TO TEXT

Calculate the break-even point in sales dollars under each approach.
Current Approach Automated Approach
Break-even point $

$

LINK TO TEXT

LINK TO TEXT

LINK TO TEXT

LINK TO TEXT

Using the current level of sales, calculate the margin of safety ratio under each approach and interpret your findings. (Round answers to 2 decimal places, e.g. 2.75%.)
Current Approach Automated Approach
Margin of safety ratio

%

%


The current approach is

lessmore

risky.

LINK TO TEXT

LINK TO TEXT

LINK TO TEXT

LINK TO TEXT

Determine the degree of operating leverage for each approach at current sales levels. (Round answers to 2 decimal places, e.g. 2.75.)
Current Approach Automated Approach
Degree of operating leverage


Calculate how much the company’s operating income would decline under either approach with a 10% decline in sales. (Round answers to 1 decimal place, e.g. 27.5%.)
In times of falling sales, the

automatedcurrent

approach would be preferred.
If there was a 10% decrease in sales, profit under the current approach would

increasedecrease

by

% compared to

% under the automated approach.

LINK TO TEXT

LINK TO TEXT

LINK TO TEXT

LINK TO TEXT

Determine at what level of sales the company’s operating income would be the same under either approach.
Level of sales $
Answer #1

A. Contribution Margin Ratio:

Current Approach Automated Approach
Contribution Margin Ratio: Contribution/Sales*100 800,000/2,000,000*100 = 40% 1,600,000/2,000,000*100 = 80%

B. Break Even Sales:

Current Approach Automated Approach

Fixed costs/Contribution Margin (%)

Fixed Costs are given in question.

Contribution Margin Ratio is as above

200,000/40*100 = 500,000 600,000/80*100= 750,000

C.Margin of Safety:

Current Approach Automated Approach

Given sales - BEP Sales

Sales is given in question

BEP Sales as computed above

2,000,000 - 500,000 = 1,500,000

1500,000/2000,000*100 = 75%

2,000,000 - 750,000 = 1250000

1250,000/2,000,000*100 = 62.5%

Interpretation: Since fixed costs element in Automated Approach is higher and the contribution ratio is also higher, the margin of safety is lesser than the current approach.

D. Degree of Operating Leverage = Contribution/Operating Income.

Current Approach Automated Approach
Degree of Operating leverage 800,000/600,000 = 1.33 1600,000/1000000 = 1.6

If sales fall by 10% =>

Current Approach Automated Approach
Sales 2,000,000 * 0.9 = 1,800,000 2,000,000*0.9 = 1,800,000
Contribution Margin 40%*1800000 = 720,000 80%*1800000 = 1440000
Fixed Costs (given) 200,000 600,000
Operating Profit 520,000 840,000

So,If there was 10% decrease in sales, the profit would be decrease by (600,000-520,000)/600,000*100 = 13.33% in current approach as compared to (1000000-840000)/1000000*100 = 16% under automated approach.

E. Indifference Point ie the point at which operating income under both level of sale are same.

Let sales be X

So,

Contribution - Fixed Cost = profit.

Hence

0.4X - 200,000 = 0.8X - 600,000

So X = 1,000,000.

Hence when sales = $ 1,000,000 operating profit is the same under both the approaches.

Please comment in case of any query regarding the solution.